E/D night 20160124

Folks: Barry, Dave, Olga, Michelle, Shane, Andrew, Marco, David G-H

Summary:

Poor weather hindered most of the useful scientific or E-night productivity, but we still managed to get one MODS science target and specphot in poor weather and do some important MODS and LBC training with Shane and Andrew during second half. During first half, Olga took some MODS2 cals and Dave tried the LUCI2 AGW check before running into MOS error (see Peter/Dave's mails)

Second Half Log:

(First half mostly closed dome testing and calibration with LUCI2 and MODS)

LBC Pointing and Collimation Test

(UT Times)

07:48 Slewing to first pointing star: WT10_361

Seeing = 1.5 - 2.0". Also cloudy, moon, and 2 degree temp difference

DOFPIA is unsurprisingly struggling. Not many pupils and no central obscuration in some.

DOFPIA failed, 'command rejected' error on LBC. Aborted with \x91S\x92, which took a bit of time.

07:59 Clear optics. Running DOFPIA again. Not using /X2 to see if we can reduce some of the brighter pupils and bg

08:06 Failed again. Same reasons. Too few pupils, and command rejected... Trying a different field, hopefully with more stars for pupils

08:09 Slewing to ACT0346

08:11 Running DOFPIA

08:15 Failed again. Looks like the red side is in focus\x85odd values on z4 (16000?)

08:18 Optics cleared again, trying DOFPIA.

08:22 Continued bad luck finding pupils and with spherical.

08:25 Slewing to ACT0282; same issues. Clouds moving in.

08:34 Giving up on pointing tests due to problems. Likely weather based and moon made it much worse. Switching to MODS, reconfiguring the telescope.

09:23 Slewing to Feige34 for BRothberg science. Andrew driving

10:14 Observing specphot.

10:00 Slewing to science target. Conditions worsen, clouds and poor seeing (>1.5") Science target is too faint to acquire

10:15 Repeating acquisition for training purposes. Shane driving.

10:32 Starting science. Could not see target through slit as weather is very poor (clouds, high background, and bad seeing, 1.3\x94), but could see bright secondary acquisition star.

10:58 The wavefront sensing looks bad. Turns out the secondary has crashed. As there is only 5 mins left in the observations we will let it finish before switching over to LBC as planned. See some faint spectra where we think target should be.

11:07 Handing over to Marco for LBC work. See his log. Very successful with some potential problems sending z11 to the telescope...

12:40 Closed again for weather (wind)

-- %USERSIG{MichelleEdwards - 2016-01-24}%

Comments:

Topic revision: r1 - 24 Jan 2016, MichelleEdwards
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback